OpenStudy (anonymous):

Just wanted to check my answer. Find the derivative of y = (e^x)(sin2x) is it = ((e^x)(cos2x))(-sin2x) ?

OpenStudy (paxpolaris):

\[y'= e^x \cdot \sin(2x)+e^x \cdot \cos(2x)\cdot2\]

OpenStudy (paxpolaris):

product rule would apply

OpenStudy (anonymous):

ahh okay, I was using the chain rule. thank you!

OpenStudy (anonymous):

sorry, if you're still around, where did the last 2 come from? after cos2x?

OpenStudy (paxpolaris):

chain rule (sin x)' = cos x (sin 2x)' = 2*sin2x

OpenStudy (paxpolaris):

sorry \[(\sin 2x)' =\cos 2x \times \left( 2x \right)'\\ = 2 \cos2x\]

OpenStudy (anonymous):

ahhh okay that makes sense, thanks again!